OpenStudy (jaylelile):

together, a washer and dryer cost $750. If the washer costs $85 more than the dryer, what does each appliance cost?

OpenStudy (jaylelile):

@Nnesha ???

OpenStudy (displayerror):

If the dryer costs $x (x dollars), how much does the washer cost?

OpenStudy (jaylelile):

I'm just having a hard time putting this into an equation...

OpenStudy (jaylelile):

x=x+85=740? I just don't see how to set this up..

OpenStudy (displayerror):

Don't worry about the equation and don't think about math. If I tell you that I have x apples and you have three more apples than I do, how many apples would you have?

OpenStudy (jaylelile):

I would have 3 more than you..

OpenStudy (anonymous):

750 - 85 = x / 2 = y

OpenStudy (anonymous):

does that help?

OpenStudy (displayerror):

If I have 2 apples, then you would have 2+3 = 5 apples. If I have x apples, then you would have x+3 apples. If the dryer costs x dollars, then how much would the washer cost?

OpenStudy (jaylelile):

x+85

OpenStudy (displayerror):

Right! Back to the apple example: If I have 2 apples and you have 2+3=5 apples, how many apples do we have altogether?

OpenStudy (jaylelile):

7 apples

OpenStudy (displayerror):

RIght! If the dryer costs x dollars and the washer costs (x+85) dollars, how much would the washer and dryer together cost?

OpenStudy (jaylelile):

whatever x is plus 85.. which is somehow 740...

OpenStudy (jaylelile):

OH WAIT IS MY EQUATION X+85=740???? IS IT THAT SIMPLE?

OpenStudy (displayerror):

Very close: The dryer costs x dollars The washer costs x+85 dollars. Adding together, we find the total price to be?

OpenStudy (jaylelile):

740..

OpenStudy (jaylelile):

I'm so confused

OpenStudy (displayerror):

I meant the equation representing the total price. In the apple example, I had 2 apples and you had 5 apples, so you said we have 7 apples. How did you get that number? Applying that to this question, if the dryer costs x dollars and the washer costs x+85 dollars, how much would the total price be? Don't think of x as a variable. What if x=0? If the dryer costs 10 dollars and the washer costs 10+85 dollars, how much would the washer+dryer cost together?

OpenStudy (jaylelile):

so just 85 dollars? I don't know how much the dryer is..

OpenStudy (displayerror):

Think of x = (one box of 3 apples). If I have x boxes and you have x boxes, how many apples do we have in total? How many boxes do we have in total?

OpenStudy (jaylelile):

6

OpenStudy (displayerror):

Right! I have x = (one box) of apples and you have x = (one box) of apples, so together we have 2x = (two boxes) of apples, which gives us a total of 6 apples. Do you see what I'm getting at here?

OpenStudy (jaylelile):

I think so......

OpenStudy (displayerror):

Don't think of x as some foreign thing. It's just a placeholder for a number. If x=1, then x+x is the same as 1+1 which equals 2. Back to your question: If the dryer costs x dollars and the washer costs x+85 dollars, then what equation gives you the total price?

OpenStudy (jaylelile):

x+85(2)=740??????

OpenStudy (displayerror):

Dryer (x) + Washer (x+85) = 740 That gets you x + x+85 = 740 Remember that you can combine your x's, just as you did when we thought of x as a box of apples.

OpenStudy (texaschic101):

this will get you the same answer.... x = washer y = dryer x + y = 740 -- washer and dryer cost 740 y = x + 85 -- dryer cost 85 more then the washer (x) and just solve by substitution but like I said, it is basically the same thing as the equation given by DisplayError ^^

OpenStudy (jaylelile):

this is what I got... the washer is $412.50 and the dryer is $327.00

OpenStudy (jaylelile):

yes? no?

OpenStudy (anonymous):

@displayError @texaschic101

OpenStudy (jaylelile):

@Hero can you check my answer for me?

OpenStudy (anonymous):

no. if you add them together it doesn't equal $750. 750 - 85 = ? /2 = ? is an equation for drier and the equation for the washer is 750 - 85 = ? /2 = ? +85

OpenStudy (anonymous):

the drier*

OpenStudy (jaylelile):

ohhhhhhh okay this time I got the washer=417.50 and the dryer=332.50

OpenStudy (anonymous):

correct.

OpenStudy (jaylelile):

yay!!! thank you so much for sticking around to help!!!

OpenStudy (anonymous):

no problem. :D